The area of a triangle is 75 square feet.If the height is 7 less than the base, find the length of the height and base of the triangle.
Answer by rothauserc(4718)

The area of a triangle is equal to 1/2 * base * height
if x is the base, then the height is x-7
75 = 1/2 * x * x-7
multiply both sides of = by 2
150 = x^2 -7x
and
x^2 -7x -150 = 0
use quadratic equation to solve for x
x = (7 + or - square root(49 - 4*1*(-150))) / 2
x = (7 + 25.48) / 2 = 16.24
x = (7 - 25.48) / 2 = -9.24
there fore we have
base is 16.24 and the height is 9.24